OpenStudy (anonymous):

find the area of the sector with 80 degree angle and radius=8

OpenStudy (anonymous):

11.164

OpenStudy (anonymous):

area of sector= angle/360 x area of circle 80/360 x \[\Pi\]r^2 =80/360 x pi x 64 =44.68units^2

OpenStudy (anonymous):

when i did the formula i got that answer too, but the answers in the back of the book say 35.45 ?

OpenStudy (anonymous):

trust urs always, textbuk may be wrong sumtyms like printing mistake etc.